### v1

# arXiv:2603.00269v1[stat.ME] 27 Feb 2026

Abstract
Linear regression estimators are known to be sensitive to outliers, and one alternative to obtain a robust and efficient estimator of the regression parameter is to model the error with Student’s $t$ distribution. In this article, we compare estimators of the degrees of freedom parameter in the $t$ distribution using frequentist and Bayesian methods, and then study properties of the corresponding estimated regression coefficient. We also include the comparison with some recommended approaches in the literature, including fixing the degrees of freedom and robust regression using the Huber loss. Our extensive simulations on both synthetic and real data demonstrate that estimating the degrees of freedom via the adjusted profile log-likelihood approach yields regression coefficient estimators with high accuracy, performing comparably to the maximum likelihood estimator where the degrees of freedom are fixed at their true values. These findings provide a detailed synthesis of $t$-based robust regression and underscore a key insight: the proper calibration of the degree of freedom is as crucial as the choice of the robust distribution itself for achieving optimal performance.
1 Introduction
Robustness is a fundamental concern in statistical inference. The presence of outliers can distort parameter estimates and degrade the reliability of inference. Some methods that have been developed for robust inference include $M$-estimation, with estimating functions designed to downweigh outliers (Huber, 1973), $L$-estimation, based on linear combinations of order statistics (Koenker and Portnoy, 1987), and $R$-estimation, based on ranks (Huber and Ronchetti, 2009, Chapter 3).
In this paper we consider robust estimation in the linear regression model
$y=X\beta+\sigma\epsilon,$ (1)
where $y=(y_{1},\dots,y_{n})$ is a vector of responses, $X$ is an $n\times p$ matrix of covariates with $i^{th}$ row $x_{i}^{T}$, and $\epsilon=(\epsilon_{1},\dots,\epsilon_{n})$ are unobserved random terms of mean zero and variance 1, independent of $y$ and $X$. Following several authors (Lange et al., 1989; Zellner, 1976; Singh, 1988; Fernández and Steel, 1999), we assume $\epsilon_{i},i=1,\dots,n$ are independent and identically distributed as Student’s $t$ with $\nu$ degrees of freedom.
Under this assumption, the likelihood function for (1) is
$L(\beta,\sigma,\nu;y)=\frac{\Gamma^{n}((\nu+1)/2)}{\sigma^{n}(\pi\nu)^{n/2}\Gamma^{n}(\nu/2)}\prod_{i=1}^{n}\left\{1+\frac{(y_{i}-x_{i}^{T}\beta)^{2}}{\nu\sigma^{2}}\right\}^{-(\nu+1)/2}.$ (2)
Whether the covariates $X$ are fixed or random, inference is based on the conditional density of $y$ given $X$, as is usual in regression problems. The estimates derived from (2) may be expected to be more robust than least-squares estimates, as the $t$-distribution has heavier tails than the normal distribution, especially for smaller values of $\nu$; see Figure 1. As $\nu\rightarrow\infty$, the $t$-density function converges to the normal density function.
The ordinary least squares estimator of $\beta$ has maximum efficiency under the normal model, but is also sensitive to outliers. Models with smaller values of $\nu$ can accommodate heavier tails, but may be inefficient if the data suggests a larger value of $\nu$. Lange et al. (1989) suggested estimating $\nu$ when $n$ is large but choosing some fixed value if $n$ is small. Liu and Rubin (1995) described an extended expectation-maximization (EM) algorithm for computing maximum likelihood estimates of the $(\beta,\sigma,\nu)$. Singh (1988) proposed a moment-based estimator of $\nu$ for $\nu>4$. Several authors have considered Bayesian inference for $\nu$, using various choices of noninformative priors, such as Jeffreys’ prior (Fonseca et al., 2008; Fernández and Steel, 1999), a reference prior (He et al., 2021) and a uniform prior (Relles and Rogers, 1977).
Figure 1: Student’s ttdistribution has heavier tails than the normal distribution, with the difference decreasing as $\nu$ increases. Asymptotically, as $\nu\rightarrow\infty$ the $t$-distribution converges to the normal distribution.
In Section 2 we compare frequentist and Bayesian approaches to the estimation of $\nu$. In Section 3 we focus on inference for $\beta$, using a two-stage procedure in which we first estimate the degrees of freedom parameter $\nu$, and then fix $\nu$ at this estimated value in the likelihood function and maximize over $\beta$ and $\sigma$. With large $p$ we find that estimating $\nu$ by maximizing the adjusted profile likelihood function gives mean-squared error for $\beta$ comparable to that from minimizing Huber loss Huber (1973). In Section 4 we consider the robustness of our proposed methods in the presence of outliers.
2 Estimating the degrees of freedom
2.1 Likelihood Approaches
The first estimator of $\nu$ we consider is the maximum likelihood estimator, $\hat{\nu}$, which maximizes the profile log-likelihood function
$\ell_{\rm p}(\nu;y)=\ell(\hat{\beta}_{\nu},\hat{\sigma}_{\nu},\nu;y),$ (3)
where $\ell(\theta)=\log L(\theta;y)$, and $(\widehat{\beta}_{\nu},\widehat{\sigma}_{\nu})=\arg\max_{(\beta,\sigma)}\ell(\beta,\sigma,\nu;y)$ are the constrained maximum likelihood estimates of $(\beta,\sigma)$. Maronna (1976) and Zellner (1976) showed that $\widehat{\beta}_{\nu}$ and $\widehat{\sigma}_{\nu}$ exist. The asymptotic properties of $\hat{\nu}$ follow from classical likelihood theory:
$\sqrt{n}\,j_{\rm p}^{1/2}(\hat{\nu})(\hat{\nu}-\nu)\xrightarrow{d}N(0,1),\quad n\rightarrow\infty,$
where $j_{\rm p}(\nu)=-\ell_{\rm p}^{\prime\prime}(\nu)$ can be computed from the full information matrix; see the supplementary material.
An adjusted profile log-likelihood function was suggested in Cox and Reid (1987) to adjust for errors in estimation of the nuisance parameters $\lambda=(\beta,\sigma$),
$\ell_{\rm adj}(\nu;y)=\ell_{\rm p}(\nu;y)-\frac{1}{2}\log|j_{\lambda\lambda}(\nu,\widehat{\lambda}_{\nu})|,$ (4)
where $j_{\lambda\lambda}(\nu,\widehat{\lambda}_{\nu})$ is the sub-matrix of the observed Fisher information matrix that corresponds to the nuisance parameters. An expression for $j_{\lambda\lambda}(\nu,\lambda)$ is given in the supplementary material. Our second likelihood-based estimate is $\hat{\nu}_{adj}$, the value that maximizes the adjusted profile log-likelihood function $\ell_{\rm adj}(\nu;y)$.
2.2 Bayesian Approaches
Jeffreys (1946) proposed non-informative priors based on the expected Fisher information matrix, which is
$I(\beta,\sigma,\nu)=\begin{pmatrix}\frac{1}{\sigma^{2}}\cdot\frac{\nu+1}{\nu+3}\sum\limits_{i=1}^{n}x_{i}x_{i}^{T}&0&0\\ 0&\frac{2n}{\sigma^{2}}\cdot\frac{\nu}{\nu+3}&-\frac{2n}{\sigma}\cdot\frac{1}{(\nu+1)(\nu+3)}\\ 0&-\frac{2n}{\sigma}\cdot\frac{1}{(\nu+1)(\nu+3)}&\frac{n}{4}\left[\psi^{\prime}\left(\frac{\nu}{2}\right)-\psi^{\prime}\left(\frac{\nu+1}{2}\right)-\frac{2(\nu+5)}{\nu(\nu+1)(\nu+3)}\right]\end{pmatrix}$. (5)
Following Fonseca et al. (2008), we consider Jeffreys’ so-called independence prior, in which the prior for the location parameter $\beta$ is treated independently of that for $(\sigma,\nu)$. Jeffreys’ rule priors are discussed in detail in Kass and Wasserman (1996). This leads to the prior
$\pi_{IJ}(\beta)\propto 1,\quad\quad\pi_{IJ}(\sigma,\nu)\propto\sigma^{-1}\left\{\frac{\nu}{\nu+3}\right\}^{\frac{1}{2}}\left\{\psi^{\prime}\left(\frac{\nu}{2}\right)-\psi^{\prime}\left(\frac{\nu+1}{2}\right)-\frac{2(\nu+3)}{\nu(\nu+1)^{2}}\right\}^{\frac{1}{2}},$ (6)
where the prior for $(\sigma,\nu)$ is the determinant of the $2\times 2$ submatrix of (5). Using these priors we maximized the posterior distribution over $(\beta,\sigma,\nu)$ and extracted the Bayesian MAP estimate $\hat{\nu}_{Bayes}$.
An alternative partially Bayes approach is to treat the profile likelihood function for $\nu$ as a true likelihood function and use the prior $\pi(\nu)\propto I_{\nu\nu}^{1/2}(\nu)$ to form a pseudo-posterior distribution. The resulting posterior distribution is proper, as proved in the supplementary material. The MAP estimate of $\nu$ obtained from maximizing this pseudo-posterior distribution is denoted by $\hat{\nu}_{Pseudo}$.
Estimates of $\nu$ from these objective functions are invariant to location-scale transformations. Specifically, for the linear model $y_{i}=x_{i}\beta+\epsilon_{i}$, transforming the response as $y^{\prime}_{i}=ay_{i}+x_{i}\vec{b}$ with $a>0$ and a constant vector $\vec{b}$ shifts the log likelihoods and log pseudo-posterior by a constant, leaving the $\nu$ optimization results unchanged. See the supplementary material for details.
Following Lange et al. (1989), we estimated the degrees of freedom $\nu$ using the reparametrization $\omega=1/\nu$. For the normal distribution, this parameterization corresponds to $\omega\rightarrow 0$ or $\nu\rightarrow\infty$. When specifying priors, we applied the change of variable: if $p(\nu)$ is the prior for $\nu$, then the prior for $\omega$ is $\pi(\omega)=p(1/\omega)\,/\omega^{2}$.
2.3 Simulations
In this section, we compare the performance of the four $\nu$-estimation approaches proposed above. We generated data from (1) with $t$ errors for $\epsilon_{i}$, with degrees of freedom $\nu=\{2,5,10\}$, $\sigma=1$, and $\beta\in\mathbb{R}^{p}$ with dimension $p\in\{1,2,5,10,20,40,60,80\}$. The covariates were generated from a multivariate standard normal distribution $\mathbf{N}_{p}(\boldsymbol{0},\mathbf{I}_{p})$. The choice of the true values of $\beta$ and $\sigma$ is arbitrary as the inference for $\nu$ is invariant to location-scale changes. The sample sizes were 300, 2500 and 4500 when $\nu=2,5,$ and $10$ (regardless of dimension $p$) respectively.
The BFGS algorithm provided in R (R Core Team, 2023) was used for all optimization steps in our simulations. This is a quasi-Newton method which iteratively approximates the Hessian of the objective function. The starting values of $(\beta,\sigma)$ for constrained maximum likelihood estimation were the ordinary least squares estimate of $\beta$ and usual residual mean square estimate of $\sigma$.
All computations were carried out in the $\omega$ parameterization, with initial guess set to the true value of $1/\nu$, but results are reported for estimation of $\nu$. The covariates were fixed in each set of simulation, to reflect modelling the conditional expected value of $y$ given $x$. We ran $500$ simulations for each combination of $\nu$ and $p$. The supplementary material provides more details about the choice of optimization details, including starting values, sample sizes and handling out-of-bound values.
p $\hat{\nu}$ $\hat{\nu}_{adj}$ $\hat{\nu}_{IJ}$ $\hat{\nu}_{Pseudo}$
1 0.3456 0.3512 0.3545 0.3577
2 0.3441 0.3523 0.3527 0.3558
5 0.3363 0.3531 0.3438 0.3466
10 0.3362 0.3655 0.3422 0.3444
20 0.3381 0.3838 0.3405 0.3411
40 0.4361 0.4488 0.4322 0.4267
60 1.2713 0.5810 1.2935 1.2486
80 1.6681 0.9010 1.6608 1.6751
Table 1: RMSE of $\hat{\nu}$ under different methods, with true $\nu=2$ and sample size $n=300$.
Table 1 presents the root mean-square-error (RMSE) averaged over the simulations, for the four estimators of $\nu$, when the true value of $\nu=2$ and the sample size $n=300$. For small values of $p$ the maximum profile likelihood estimator $\hat{\nu}$ has the smallest RMSE, but all four methods are comparable. For $p=60$ or $80$, the adjusted profile likelihood estimator $\hat{\nu}_{adj}$ has the smallest RMSE, and seems to adapt well to high-dimensional settings. The Bayesian estimators have similar RMSE when compared to each other and to the likelihood estimators. Other choices of $\nu$, $n$, and $p$, reported in the supplementary material, give similar conclusions. If $p$ is not large compared to $n$, we recommend the $\hat{\nu}$ and for larger $p$, $\hat{\nu}_{adj}$.
3 Estimating the regression
3.1 Simulation Settings
In linear regression, estimating the coefficient is the primary objective. To evaluate the effect of $\nu$-estimation in supporting coefficient estimation, we designed a simulation study using the stack loss data of Brownlee (1965). It has 21 responses ($Y$) in an industrial process, along with three explanatory variables: air flow ($X_{1}$), temperature ($X_{2}$), and acid concentration ($X_{3}$). We conducted simulations according to the following three data generation methods:
  1. 1.
    In the first method, we used the original design matrix $X$ from the stackloss dataset, which contains $n=21$ observations and $p=4$ predictors (including an intercept). The least-squares estimates of $\beta$ and $\sigma$ for this dataset were treated as the true values. We generated response values according to
    $y_{i}=x_{i}^{T}\beta+\epsilon_{i},\,i=1,\dots,21,$ (7)
    with $X$ from the original data, and $t_{2}$ errors $\epsilon_{i}$ .
  2. 2.
    In the second method, we expanded the data to $p=\{40,80,120\}$ and $n=210$ to study the performance in a higher-dimensional regime with a larger sample size. The first four predictors (including the intercept) were based on the original stackloss predictors, in which we retained the original 21 observations, and generated an additional 189 observations with 3 predictors from a multivariate normal distribution with the same empirical mean vector and covariance matrix obtained from the original design matrix. All the remaining p4p-4 predictors were generated independently from a standard normal distribution. The corresponding true $\beta$ vector had the first four coefficients being the OLS estimates from the original data, and the remaining coefficients set as zero.
  3. 3.
    In the third method, we generated the data using (7), with $\epsilon_{i}$ following either $t_{2}$ or $N(0,1)$ to understand the effect of error distribution. Data was simulated with these $(n,p)$ settings: $(20,4)$, $(500,4)$, and $(100,50)$.
We considered three approaches for the estimation of $\nu$: profile likelihood $\hat{\nu}_{p}$, adjusted profile likelihood $\hat{\nu}_{adj}$, and Jeffreys $\hat{\nu}_{Bayes}$. We also considered some fixed value other than the true $\nu=2$. For Bayesian methods, we only considered the Jeffreys approach because both Bayesian methods give very similar $\nu$-estimation in Section 2, and full independence Jeffreys is easier to implement, as it does not first construct the profile likelihood function. In each simulation, we estimated or fixed $\nu$ from the four approaches, then obtained $(\widehat{\beta}_{\nu},\widehat{\sigma}_{\nu})$ by optimizing the Student’s $t$ likelihood function. We also considered the ordinary least squares (OLS) and Huber regression estimators (Huber, 1973) of $\beta$ for comparison. For Huber’s method, the tuning parameter was automatically selected using the rlmDataDriven R package (Wang et al., 2019). This procedure was repeated over 500 simulations, again keeping the design matrix $X$ fixed while varying only the error term.
OLS imposes the largest penalty on the outliers compared to other methods. $t_{1}$ imposes lighter penalty than Huber regression with tuning parameter $c=1.5$). As $\nu$ increases, the loss function behaves more similarly as OLS loss function. See Figure 2.
Figure 2: Loss function of OLS vs Huber vs t distribution
The simulation RMSE for $\hat{\beta}$ was computed as the average of the square root MSE over the scalar, ignoring the intercept
$RMSE(\hat{\beta})=\frac{1}{500}\sum_{k=1}^{500}\sqrt{\frac{1}{p-1}\sum_{j=2}^{p}(\beta_{jk}-\hat{\beta}_{jk})^{2}}.$ (8)
where $\hat{\beta}_{jk}$ is the $k$th simulation’s estimate of $j$th coefficient.
3.2 Results
Figure 3: Overall RMSE of $\hat{\beta}$ using stackloss data (data generation method 1)
(a) $n=21,p=4$
(b) $n=210,p=40$
(c) $n=210,p=80$
(d) $n=210,p=120$
The plots compare fixed $\nu$ approaches (in blue) for $\nu\in\{1,\dots,10\}\cup\{20,30\}$ and other traditional estimators (OLS and Huber) with the proposed $\nu$-estimation approaches (in red).
From Figure 3(a), we observe that the lowest overall RMSE is achieved when fixing $\nu=3$, with the RMSE at the true value $\nu=2$ is only slightly higher, indicating that the method is not highly sensitive to small misspecifications of $\nu$ around the true value. The slightly better performance at $\nu=3$ compared to the true $\nu=2$ is likely due to sampling variability in the simulation. As $n$ increases, the true $\nu$ would result in the lowest RMSE. For higher values of $\nu$ (e.g., $\nu=20$ or $\nu=30$), the RMSE for the estimate of $\beta$ increases notably. The three $\nu$-estimation approaches are comparable to the best fixed-$\nu$ results, and all outperform ordinary least squares. Estimation of $\beta$ using $\hat{\nu}_{adj}$ gives the smallest RMSE among $\nu$ estimated methods, which is comparable to RMSE of fixed $\nu=6$. Huber regression is slightly better than $\hat{\nu}_{adj}$ approach.
According to Figures 3(b) and 3(c), when $p=40,80$ ($p/n\approx 0.190,0.381$ respectively), the lowest overall RMSE is achieved by fixing $\nu$ at $3$ or $4$, and the RMSE increases for $\nu>4$. The $\hat{\nu}_{adj}$ approach outperforms Huber regression and gives the lowest RMSE among all $\nu$ estimated methods, which is comparable to fixing at the true $\nu$. In Figure 3(d), as we further increase $p$ to $120$ ($p/n\approx 0.571$), the lowest overall RMSE is achieved at larger fixed $\nu=7$. High dimensionality of $\beta$ gives the model enough flexibility to absorb extreme values into the fitted coefficients, shrinking the residuals relative to the true $t_{2}$ errors. With these outliers muted, the likelihood favours a larger degrees of freedom $\nu$, similar to selecting a less robust (steeper) loss function in Figure2, since the data no longer exhibit the heavy-tailed behaviour that would justify a small $\nu$. Among the $\nu$-estimation approaches, $\hat{\nu}_{adj}$ approach performs the best, its RMSE value is relatively large compared to the lowest RMSE that could be achieved by traditional methods. In particular, Huber regression performs better than the $\hat{\nu}_{adj}$ approach.
Figure 4: Overall RMSE of $\hat{\beta}$ in simulated $t$-error data (data generation method 2)
(a) $n=20,p=4$
(b) $n=500,p=4$
(c) $n=100,p=50$
The plots compare non $\nu$-estimation approaches (in blue) across various fixed values of $\nu$ and other traditional estimators (OLS and Huber) with the proposed $\nu$-estimation approaches (in red) in terms of overall RMSE.
According to Figures 4(a) and 4(b), the trend in overall RMSE using the simulated data with t error is similar to that of the stack loss data when $p/n<0.5$, where RMSE tends to increase as the fixed $\nu$ shifts away from the true value. The $\hat{\nu}_{adj}$ approach results in a lower RMSE than Huber regression. According to Figure 4(c), when $p/n=0.5$, we observe a similar RMSE trend as the stackloss $n=210,p=120$ simulations. The lowest RMSE is achieved at fixed $\nu=6$, which is not close to true $\nu=2$. Huber regression performs better than the $\hat{\nu}_{adj}$ approach, although both approaches give relatively large RMSEs compared to the smallest RMSE given by fixing $\nu=6$.
Figure 5: Overall RMSE of $\hat{\beta}$ using simulated normal-error data (data generation method 3)
(a) $n=20,p=4$
(b) $n=500,p=4$
(c) $n=100,p=50$
The plot compares non $\nu$-estimation approaches (in blue) across various fixed values of $\nu$ and other traditional estimators (OLS and Huber) with the proposed $\nu$-estimation approaches (in red) in terms of overall RMSE.
With standard normal errors (Figures 5(a) and 5(b)), OLS performs the best in both $n=20$ and $n=500$ cases. All $\nu$-estimation approaches outperform Huber regression with $\hat{\nu}_{adj}$ approach coming closest to OLS among the $\nu$-estimation approaches. In the high-dimensional case $n=100;p=50$, Figure 5(c) shows that OLS again performs the best and only $\hat{\nu}_{adj}$ surpasses Huber regression among the $\nu$-estimation approaches. Overall, estimating $\nu$ is beneficial when when $p/n$ is not too large, as it would give promising results regardless the underlying error distribution. For errors with light-tail distribution, adjusted profile likelihood approach is comparable to the OLS, and for heavy tail distribution, it outperforms Huber when $p/n$ ratio is below 0.5. Although Lange et al. (1989) recommends fixing $\nu$ for small sample size $n$, the adjusted profile likelihood approach yields performance comparable to the best fixed-$\nu$ choices across error settings, supporting its use for estimating $\nu$.
4 Robustness to outliers
4.1 Simulation Settings
To investigate $t$ regression on data with outliers, we considered a linear model with true $\beta=(1,5,10)$ and $n=100$. Additionally, we also considered data with outliers under high dimensional setting, generated using a linear model with $p=80$, true $\beta=\boldsymbol{0}_{p}$, and $n=300$. The covariates and errors were generated from the standard normal distribution, with contamination (i) $N(0,9)$, (ii) $\chi^{2}(4)-4$, (iii) $t(2)$, or (iv) two-point contamination in which any response value may become 5-5 or 5 with a probability $\lambda/2$ each. The contamination rates considered are $\lambda=$ 10%, 20%, and 30%.
4.2 Results when $p=3$
Figure 6: Overall RMSE of $\hat{\beta}$ with 10%, 20%, and 30% $N(0,9)$ contaminated errors when $p=3$
Refer to caption Refer to caption Refer to caption
When the data is contaminated with $N(0,9)$ errors (see Figure 6) , although the $\nu$-estimation approaches do not necessarily give the lowest overall RMSE, they are able to give relatively low RMSE compared to OLS and Huber regression. As the contamination rate increases, the $\nu$-estimation approaches outperform most of the fixed $\nu$ approach estimations as they are successful in estimating $\nu$ close to truth. Similar pattern is shown in data with two-point contamination.
Figure 7: Overall RMSE of $\hat{\beta}$ with 10% 20%, and 30% $\chi^{2}(4)-4$ contaminated errors when $p=3$
Refer to caption Refer to caption Refer to caption
[…]
4.3 Results when $p=80$
In high dimension setting, when the data is contaminated with $N(0,9)$ (see Figure 8), $\chi^{2}(4)-4$, or $t(2)$ errors, although $\hat{\nu}_{adj}$ approach does not necessarily give the lowest overall RMSE, it is able to give relatively lower RMSE compared to OLS and Huber regression. This may be explained by the fact that $\hat{\nu}_{adj}$ are closer the the truth compared to other $\nu$ estimation methods (and fixing large $\nu$) in high dimension setting.
Figure 8: Overall RMSE of $\hat{\beta}$ with 10%, 20%, and 30% $N(0,9)$ contaminated errors when $p=80$
Refer to caption Refer to caption Refer to caption
Figure 9: Overall RMSE of $\hat{\beta}$ with 10%, 20%, and 30% two-point contaminated errors when $p=80$
Refer to caption Refer to caption Refer to caption
For data contaminated with two-point errors (see Figure 9), $\hat{\nu}_{adj}$ approach outperforms Huber and OLS when the contamination rate is low (e.g. 10%). As the contamination rate increases, Huber regression performs better as $\hat{\nu}_{adj}$ are overestimated, and hence the corresponding $\beta$ estimation performance is similar to fixing $\nu=30$. Nonetheless, the $\nu$-estimation approaches outperform OLS when the contamination rate is 20%. As the contamination rate increases to 30%, all $\nu$-estimation approaches perform as bad as most fixed-$\nu$ approaches and OLS. This may be caused by high contamination of outliers with a ”uniform” pattern in the data that t distribution fails to capture.
These results collectively suggest that the $\nu$-estimation approaches, and $\hat{\nu}_{adj}$ in some high dimension settings, offer a robust performance across various contamination types and levels, often outperforming classical regression methods.
5 Conclusions and Discussion
In this paper, we studied robust regression with Student’s t error and addressed the commonly asked question on how to choose the degree of freedom parameter in Student’s t distribution in practice. We compared 4 frequentist and Bayesian methods in estimating the degree of freedom $\nu$, and studied how these choices propagate to estimation of the regression coefficients $\beta$ in uncontaminated and contaminated data. In terms of $\beta$ estimation performance, we suggest estimating $\nu$ with adjusted profile likelihood when the $p/n$ ratio is not too large as it is the most robust in both heavy-tailed and light-tailed distribution, as well as in contaminated data and large $p$ (given moderate $n$) settings. However, to ensure stable $\nu$ and $\beta$ estimations, our proposed $\nu$-estimation methods require $p$ to be not too large relative to $n$. This is because high dimensionality may lead to larger residuals magnitude and tend to overestimate $\nu$. Therefore, when $p/n$ ratio is very large (it would be better if we were able to quantify this), optimizing the likelihood by fixing $\nu$ at some large values may be more efficient.
Appendix A Supplementary Material
Appendix B Derivation of observed Fisher information for $\beta$ and $\sigma$
$\ell(\beta,\sigma,\nu;y,x)$ $=c(\nu)-\frac{\nu+1}{2}\sum_{i=1}^{n}\log\{\nu+(\frac{y_{i}-x_{i}^{T}\beta}{\sigma})^{2}\}-n\log(\sigma)$
The first derivative with respect to $\beta$ and $\sigma$ are:
$\frac{\partial\ell}{\partial\beta}$ $=(\nu+1)\sum_{i=1}^{n}x_{i}^{T}\frac{y_{i}-x_{i}^{T}\beta}{\nu\sigma^{2}+(y_{i}-x_{i}^{T}\beta)^{2}}\quad\quad\frac{\partial l}{\partial\sigma}=-\frac{n}{\sigma}+\frac{\nu+1}{\sigma}\sum_{i=1}^{n}\frac{(y_{i}-x_{i}^{T}\beta)^{2}}{\nu\sigma^{2}+(y_{i}-x_{i}^{T}\beta)^{2}}.$
Define $r_{i}:=y_{i}-x_{i}^{T}\beta$, then $z_{i}:=r_{i}/\sigma$, and also define
$g_{i}(\sigma):=\frac{r_{i}^{2}}{\nu\sigma^{2}+r_{i}^{2}}\quad\quad g_{i}^{\prime}(\sigma)=\frac{-2\sigma\nu r_{i}^{2}}{(\sigma^{2}\nu+r_{i}^{2})^{2}}.$
The observed Fisher information matrix entries for $\theta=(\beta,\sigma)$ is
$\frac{\partial^{2}\ell}{\partial\sigma^{2}}$ $=\frac{n}{\sigma^{2}}-(\nu+1)\sum_{i=1}^{n}[\frac{z_{i}^{2}}{\sigma^{2}\nu+r_{i}^{2}}+\frac{2\nu r_{i}^{2}}{(\sigma^{2}\nu+r_{i}^{2})^{2}}]\quad\frac{\partial^{2}\ell}{\partial\beta^{2}}=\frac{\nu+1}{\sigma^{2}}\sum_{i=1}^{n}\frac{-x_{i}x_{i}^{T}}{\nu+z_{i}^{2}}+\frac{2z_{i}^{2}x_{i}^{T}}{(\nu+z_{i}^{2})^{2}}$
$\frac{\partial^{2}\ell}{\partial\beta\partial\sigma}$ $=-\frac{(\nu+1)2\nu}{\sigma^{2}}\sum_{i=1}^{n}\frac{x_{i}^{T}z_{i}}{(\nu+z_{i}^{2})^{2}}$
Appendix C Proof of Invariance Property
[…]
C.1 Profile likelihood
The profile log-likelihood function is
$\ell^{P}(\nu;y,x)$ $=C(\nu)-\frac{\nu+1}{2}\sum_{i=1}^{n}\log\{\nu+z_{i}^{2}\}-n\log(\hat{\sigma}_{\nu}).$
Then, for a new response $y^{\prime}$,
$l^{P}(\nu;y^{\prime},x)$ $=C(\nu)-\frac{\nu+1}{2}\sum_{i=1}^{n}log\{\nu+z_{i}^{2}\}-n\log(a\hat{\sigma}_{\nu})=l^{P}(\nu;y,x)-n\log(a),$
so the log profile likelihood function is shifted by $n\log(a)$.
C.2 Adjusted Profile likelihood
By the Schur complement formula,
[…]
The submatrix $j_{11}$ is a $p\times p$ matrix:
$j_{11}$ $=(\nu+1)\sum_{i=1}^{n}x_{i}x_{i}^{T}\frac{\nu\sigma^{2}-r_{i}^{2}}{(\nu\sigma^{2}+r_{i}^{2})^{2}}=\frac{1}{\sigma^{2}}\times M.$
The $p\times p$ matrix $M$ is independent of $\sigma$. Therefore, $|j_{11}|=\frac{1}{\sigma^{2p}}|M|\propto\frac{1}{\sigma^{2}}$. The component $j_{12}$ is a $p\times 1$ vector:
$j_{12}=\frac{(\nu+1)2\nu}{\sigma^{2}}\sum_{i=1}^{n}\frac{x_{i}^{T}z_{i}}{(\nu+z_{i}^{2})^{2}}\propto\frac{1}{\sigma^{2}}\text{ and }j_{22}=-\frac{n}{\sigma^{2}}+\frac{\nu+1}{\sigma^{2}}\sum_{i=1}^{n}\frac{z_{i}^{2}}{\nu+z_{i}^{2}}+\frac{2\nu z_{i}^{2}}{(\nu+z_{i}^{2})^{2}}\propto\frac{1}{\sigma^{2}}.$
Therefore, we have
$|j_{\lambda\lambda}(\nu,\hat{\lambda}_{\nu})|\propto\sigma^{-2p-2}\quad\text{ and }\quad-\frac{1}{2}\log|j_{\lambda\lambda}(\nu,\hat{\lambda_{\nu}})|=K+(p+1)\log(\sigma),$
where $K$ is independent of $\sigma$. This shows that the profile log-likelihood function is shifted by $n\log(a)+(p+1)\log(a)$ under regression-scale transformation.
C.3 Full Independence Jeffreys
For Bayesian methods, the posterior is proportional to its joint density, which consists of a likelihood part and the prior part. For log-likelihood, similar to profile log-likelihood, it is shifted by $n\log(a)$. Meanwhile, the log $\sigma$ prior would be shifted by $\log(a)$. So, the log Full Jeffrey’s joint density would be shifted by $(n+1)\log(a)$.
C.4 Marginal Jeffrey’s
For both Marginal Jeffrey’s methods (Independence and $\nu$ block), since we do not impose priors on $\sigma$, only the profile likelihood part consists of $\sigma$. Therefore, they would both be shifted the same as the profile log-likelihood.
Appendix D More details on $\nu$-optimization
D.1 Out of bound values
Using the $\omega=1/\nu$ parametrization, for any $\omega<0$, we set the functions to return infinity to avoid returning negative $\hat{\nu}$; when $\omega=0$, we treated the error distribution as normal when deriving the likelihood part of the objective functions. Finally, we inverted $\hat{\omega}$ to obtain $\hat{\nu}$.
D.2 Choice of sample sizes and starting values
Fonseca et al. (2008) noted that, as a function of $\nu$, the profile log-likelihood function will increase without bound when $\sum_{i=1}^{n}(\hat{z_{i}}^{2}-1)^{2}<2n$ where $\hat{z_{i}}=(y_{i}-x_{i}^{T}\hat{\beta)}/\hat{\sigma}$, $\hat{\beta},\hat{\sigma}$ are the OLS estimators. Hence, to ensure that the optimization algorithm would converge, we first searched for large enough $n$ such that Fonseca et al. (2008)’s condition is violated. Next, we searched for $n$ such that all 5 objective functions do not have flatness issue by manually plotting out the functions for 20 simulated datasets. When we ran simulations using smaller $n$, for instance, $\nu=2$; $p=40$ using $n=100$, around 3.33% of the simulations had flatness issue and over 20% $\hat{\nu}_{adj}$ from the remaining successful converged simulations were greater than $10$, which are bad estimates. When $p$ was further increased to $60$, the flatness issue appeared in 55% of the simulations, and over 18% $\hat{\nu}_{adj}$ from the remaining successful converged simulations were greater than $10$.
Figure 10: Boxplots $p=40$ and $p=60$ with $n=100$, $\nu=2$.
(a) (a) $p=40$
(b) (b) $p=60$
These boxplots displays the spread of $\hat{\nu}$ in successfully converged simulations under different approaches in (a) $p=40$ and (b) $p=60$ with $n=100$, $\nu=2$ and initial value set at truth settings.
Note that the choice of initial guess $\omega_{\text{init}}$ only determines if the optimization algorithm would run or not. In cases of flatness issues, for some initial value, even though the algorithm may still run and claim to converge successfully, their estimated values are not necessarily reliable (i.e. the very overestimated $\hat{\nu}_{adj}$ shown above). Hence, in our simulation analysis, we try to avoid these cases, by setting sufficiently large $n$.
When there are no flatness issue (i.e., the simulation settings we consider), setting the initial value at true $\omega$ that was used to generate the data optimizes the optimization time since we expect the returned $\hat{\omega}$ to be close to the truth. In our simulation settings, using non-truth value as initial guess would still yield to stable estimates and RMSE results under the adjusted likelihood approach. The other 4 approaches tend to underestimate $\nu$, such phenomenon becomes more obvious as we set $\omega_{\text{init}}$ to be lower than the truth. For the purpose to compare the performances between the 5 approaches, the overall RMSE conclusion were consistent with results obtained from setting the initial guess at truth, in which we concluded that the adjusted likelihood approach resulted in lowest RMSE. For example, when we ran simulations on $n=300$; $\nu=2$; $p=60$ using the $1/5$ as initial guess of $\omega$, RMSE of $\hat{\nu}_{adj}$ is $0.5824363$, which is very close to that from setting initial guess at $1/2$ (i.e., truth). The other approaches yield RMSE at around 1.361.391.36-1.39, which are slightly inflated from setting initial guess at truth which gave 1.251.291.25-1.29 RMSE. If the initial guess of $\omega$ is set at $1/10$, RMSE of $\hat{\nu}_{adj}$ is $0.5860144$ and the other approaches yield RMSE at around 1.441.491.44-1.49. The boxplots below display the spread of $\hat{\nu}$ in each setting.
Figure 11: Boxplots of $\hat{\nu}$ under different $\omega_{\text{init}}$ values
(a) (a) $\omega_{\text{init}}=1/2$
(b) (b) $\omega_{\text{init}}=1/5$
(c) (c) $\omega_{\text{init}}=1/10$
These 3 boxplots of $\hat{\nu}$ optimized with different $\omega_{\text{init}}$ values with (a) at truth, (b) at 1/5, and (c) at 1/10. The red line indicates the true $\nu$ value.
Appendix E Extra Results of $\nu$ estimation simulations
Method p = 1 p = 2 p = 5 p = 10 p = 20 p = 40 p = 60 p = 80
Profile 0.5405 0.5403 0.5388 0.5390 0.5358 0.5371 0.5372 0.5430
Adj Profile 0.5412 0.5416 0.5418 0.5450 0.5465 0.5531 0.5566 0.5556
Jeffrey’s Full 0.5440 0.5437 0.5420 0.5420 0.5383 0.5387 0.5380 0.5422
Marg Jeff ($\nu$ block) 0.5449 0.5447 0.5429 0.5428 0.5390 0.5391 0.5377 0.5422
Table 2: RMSE of $\hat{\nu}$ under different methods, with true $\nu=5$ ($n=2500$)
Method p = 1 p = 2 p = 5 p = 10 p = 20 p = 40 p = 60 p = 80
Profile 1.5032 1.5008 1.4965 1.4946 1.4863 1.4750 1.4883 1.4670
Adj Profile 1.5020 1.5015 1.5005 1.5032 1.5014 1.5123 1.5387 1.5369
Jeffrey’s Full 1.5144 1.5118 1.5072 1.5039 1.4962 1.4870 1.4959 1.4739
Marg Jeff ($\nu$ block) 1.5170 1.5146 1.5100 1.5083 1.4995 1.4897 1.4983 1.4761
Table 3: RMSE of $\hat{\nu}$ under different methods, with true $\nu=10$ ($n=4500$)
Method p = 1 p = 2 p = 5 p = 10 p = 20 p = 40 p = 60 p = 80
Profile 0.07597 0.06875 0.04292 0.01031 -0.06994 -0.28458 -1.15143 -1.65848
Adj Profile 0.08858 0.08832 0.08517 0.09308 0.10556 0.16210 0.31432 0.62606
Jeffrey’s Full 0.09234 0.08502 0.05897 0.02606 -0.05509 -0.27202 -1.17377 -1.65165
Marg Jeff ($\nu$ block) 0.09865 0.09138 0.06549 0.03291 -0.04737 -0.26154 -1.11832 -1.66641
Table 4: Bias of $\hat{\nu}$ under different methods, with with true $\nu=2$ ($n=300$)
Method p = 1 p = 2 p = 5 p = 10 p = 20 p = 40 p = 60 p = 80
Profile 0.06045 0.05722 0.04814 0.03471 0.00962 -0.04086 -0.09318 -0.1539
Adj Profile 0.06375 0.06302 0.06159 0.06106 0.06178 0.06391 0.06639 0.06302
Jeffrey’s Full 0.07084 0.06759 0.05850 0.04505 0.01980 -0.03078 -0.08341 -0.14428
Marg Jeff ($\nu$ block) 0.07379 0.07055 0.06145 0.04799 0.02279 -0.02777 -0.08009 -0.14089
Table 5: Bias of $\hat{\nu}$ under different methods, with with true $\nu=5$ ($n=2500$)
Method p = 1 p = 2 p = 5 p = 10 p = 20 p = 40 p = 60 p = 80
Profile 0.23342 0.22895 0.21841 0.20142 0.16987 0.10004 0.03927 -0.03360
Adj Profile 0.23480 0.23370 0.23339 0.23255 0.23607 0.23669 0.23806 0.24420
Jeffrey’s Full 0.25715 0.25238 0.24176 0.22428 0.19364 0.12363 0.06218 -0.01048
Marg Jeff ($\nu$ block) 0.26313 0.25850 0.24798 0.23139 0.19978 0.12882 0.06858 -0.004233
Table 6: Bias of $\hat{\nu}$ under different methods, with with true $\nu=10$ ($n=4500$)
Method p = 1 p = 2 p = 5 p = 10 p = 20 p = 40 p = 60 p = 80
Profile 0.33747 0.33753 0.33392 0.33635 0.33112 0.33073 0.53940 0.17883
Adj Profile 0.34017 0.34136 0.34301 0.35376 0.36940 0.41888 0.48912 0.64864
Jeffrey’s Full 0.34256 0.34263 0.33903 0.34154 0.33632 0.33615 0.54398 0.17451
Marg Jeff ($\nu$ block) 0.34414 0.34423 0.34066 0.34318 0.33809 0.33745 0.55581 0.17077
Table 7: Standard error of $\hat{\nu}$ under different methods, with with true $\nu=2$ ($n=300$)
Method p = 1 p = 2 p = 5 p = 10 p = 20 p = 40 p = 60 p = 80
Profile 0.53769 0.53784 0.53716 0.53844 0.53626 0.53604 0.52960 0.52134
Adj Profile 0.53792 0.53841 0.53882 0.54206 0.54355 0.54997 0.55319 0.55259
Jeffrey’s Full 0.53989 0.54003 0.53935 0.54065 0.53851 0.53833 0.53207 0.52316
Marg Jeff ($\nu$ block) 0.54046 0.54061 0.53993 0.54123 0.53910 0.53890 0.53228 0.52410
Table 8: Standard error of $\hat{\nu}$ under different methods, with with true $\nu=5$ ($n=2500$)
Method p = 1 p = 2 p = 5 p = 10 p = 20 p = 40 p = 60 p = 80
Profile 1.48642 1.48470 1.48199 1.48245 1.47807 1.47308 1.48930 1.46806
Adj Profile 1.48507 1.48467 1.48374 1.48662 1.48422 1.49514 1.52169 1.51890
Jeffrey’s Full 1.49389 1.49207 1.48920 1.48859 1.48506 1.48332 1.49608 1.47536
Marg Jeff ($\nu$ block) 1.49545 1.49387 1.49100 1.49196 1.48763 1.48556 1.49827 1.47753
Table 9: Standard error of $\hat{\nu}$ under different methods, with with true $\nu=10$ ($n=4500$)
Appendix F Order of Bias and Variance for $\hat{\nu}$
Theorem 1.
For the profile log-likelihood estimation of $\hat{\nu}$, both the variance and the bias has order $\mathcal{O}(n^{-1})$.
Proof.
From Reid (2013), asymptotically, the profile log-likelihood is
$j_{p}^{1/2}(\hat{\nu})\,(\hat{\nu}-\nu)\;\xrightarrow{\mathcal{L}}\;\mathcal{N}(0,I),$ (9)
[…]
$\frac{\partial^{2}\ell_{p}(\nu,\hat{\lambda}_{\nu};y,x)}{\partial\nu^{2}}$ $=\ell_{\nu\nu}+\ell_{\nu\lambda}\frac{\partial\hat{\lambda}_{\nu}}{\partial\nu}$
Define $f(\nu,\hat{\lambda}_{\nu})=\partial\ell(\nu,\hat{\lambda}_{\nu};y,x)/\partial\hat{\lambda}_{\nu}=0:=\ell_{\lambda}$. If $\partial f(\nu,\hat{\lambda}_{\nu})/\partial\hat{\lambda}_{\nu}$ is invertiable, taking the derivative and set $\partial f(\nu,\hat{\lambda}_{\nu})/\partial\nu=0$
$\frac{\partial\hat{\lambda}_{\nu}}{\partial\nu}=-\left(\frac{\partial f(\nu,\hat{\lambda}_{\nu})}{\partial\hat{\lambda}_{\nu}}\right)^{-1}\frac{\partial f(\nu,\hat{\lambda}_{\nu})}{\partial\hat{\lambda}_{\nu}}.$
Therefore, we have
$j_{p}(\hat{\nu})=-\ell_{\nu\nu}+\ell_{\nu\lambda}(\ell_{\lambda\lambda})^{-1}\ell_{\lambda\nu}.$
The second derivative of the log-likelihood with respect to $\nu$ is
$\ell_{\nu\nu}=\frac{n}{4}\left\{\psi^{\prime}\left(\frac{\nu+1}{2}\right)-\psi^{\prime}\left(\frac{\nu}{2}\right)\right\}+\frac{1}{2}\sum_{i=1}^{n}\left\{\frac{1}{\nu}-\frac{1}{\nu+\hat{z}_{i\nu}^{2}}-\frac{\hat{z}_{i\nu}^{2}-1}{(\nu+\hat{z}_{i\nu}^{2})^{2}}\right\}=\mathcal{O}(n).$
And the second derivative of the log-likelihood with respect to $\nu$ and $\lambda$ is
$\ell_{\nu\beta}=\sum_{i=1}^{n}-\frac{1}{\hat{\sigma}_{\nu}^{2}}\frac{(\hat{z}_{i\nu}^{2}-1)(y_{i}-x_{i}^{T}\beta)x_{i}}{(\nu+\hat{z}_{i\nu}^{2})^{2}}=\mathcal{O}(n)\quad\ell_{\nu\sigma}=\frac{1}{\hat{\sigma_{\nu}}^{3}}\sum_{i=1}^{n}\frac{(\hat{z}_{i\nu}^{2}-1)(y_{i}-x_{i}^{T}\beta)^{2}}{(\nu+\hat{z}_{i\nu}^{2})^{2}}=\mathcal{O}(n)$
The term $\ell_{\lambda\lambda}$ is the observed Fisher information matrix for the adjusted profile likelihood, $j_{\lambda\lambda}$ (LABEL:eq:obsfisher), which has order of $\mathcal{O}(n)$. Therefore, $j_{p}(\hat{\nu})=\mathcal{O}(n)$,so we have the $Var(\hat{\nu})=\mathcal{O}(n^{-1})$.
Next, we analyze the bias part, we assume the maximum likelihood estimate of $\beta,\sigma$ and $\nu$ exists. The maximum profile likelihood estimates coincides with the maximum likelihood estimates, so it suffices to show the order of bias for the maximum likelihood estimate of $\nu$. We follow the proof by Cox and Snell (1968). The maximum likelihood equation $\ell^{\prime}(\hat{\nu},\hat{\beta},\hat{\sigma};y,x)=0$ to the first order for $\nu$ is
$\ell^{\prime}(\nu)+(\hat{\nu}-\nu)\ell^{\prime\prime}(\nu)=0.$
Let
$U^{i}=\frac{\partial\log\ell(\nu,\hat{\beta},\hat{\sigma};y,x)}{\partial\nu},V^{i}=\frac{\partial\log\ell(\nu,\hat{\beta},\hat{\sigma};y,x)}{\partial\nu^{2}},$
it would be exactly $\frac{\partial\ell_{p}(\nu;y,x)}{\partial\nu}$ and $\ell_{\nu\nu}$in the previous part with the $\hat{\beta}_{\nu},\hat{\sigma}_{\nu}$ replaced by the maximum likelihood estimate, respectively. Then, we replace $\ell^{\prime\prime}(\nu)$ by its expectation
$I:=\sum_{i=1}^{n}E(-V^{i})=\frac{n}{4}\left\{\psi^{\prime}\left(\frac{\nu}{2}\right)-\psi^{\prime}\left(\frac{\nu+1}{2}\right)-\frac{2(\nu+5)}{\nu(\nu+1)(\nu+3)}\right\}$
where I is the total information in the sample, shown in Fonseca et al. (2008). Then, we have the standard first-order expressions
$\hat{\nu}-\nu=\frac{\sum_{i=1}^{n}U^{i}}{I},\quad Var(\hat{\nu})=\frac{1}{I}.$
To obtain a more refined answer, we replace it by the second-order equation
$\ell^{\prime}(\nu)+(\hat{\nu}-\nu)\ell^{\prime\prime}(\nu)+\frac{1}{2}(\hat{\nu}-\nu)^{2}\ell^{\prime\prime\prime}(\nu)=0,$
and take expectation we have
$E(\hat{\nu}-\nu)E(\ell^{\prime\prime}(\nu))+cov(\hat{\nu}-\nu,\ell^{\prime\prime}(\nu))+\frac{1}{2}E(\hat{\nu}-\nu)^{2}E(\ell^{\prime\prime\prime}(\nu))+cov(0.5(\hat{\nu}-\nu),\ell^{\prime\prime\prime}(\nu))=0$
Approximately,
$cov(\hat{\nu}-\nu,\ell^{\prime\prime}(\nu))=cov\left(\frac{\sum_{i=1}^{n}U^{i}}{I},V\right)=\frac{1}{I}cov\left(\sum_{i=1}^{n}U^{i},\sum_{i=1}^{n}V^{i}\right)=\frac{J}{I},$
where $J=\sum_{i=1}^{n}E(U^{i}V^{i})$, by independence between observation and 0 expectation of score. And if
$W^{i}=\frac{\partial^{3}\ell(\nu)}{\partial\nu^{3}}=\frac{n}{8}\left\{\psi^{\prime\prime}\left(\frac{\nu+1}{2}\right)-\psi^{\prime\prime}\left(\frac{\nu}{2}\right)\right\}+\frac{1}{2}\sum_{i=1}^{n}\left\{-\frac{1}{\nu^{2}}+\frac{1}{(\nu+\hat{z}_{i\nu}^{2})^{2}}+\frac{\hat{z}_{i\nu}^{2}-1}{(\nu+\hat{z}_{i\nu}^{2})^{3}}\right\},$
we take the expectation for the sum of $W^{i}$
$K$ $:=E(\sum_{i=1}^{n}W^{i})=E(\ell^{\prime\prime\prime}(\nu))=\frac{n}{8}\left\{\psi^{\prime\prime}\left(\frac{\nu+1}{2}\right)-\psi^{\prime\prime}\left(\frac{\nu}{2}\right)\right\}-\frac{n}{2}\frac{2\nu^{2}+21\nu+31}{\nu^{2}(\nu+1)(\nu+3)(\nu+5)}.$
$I,J,K$ refer to a total over the sample and are of order n.
$cov(0.5(\hat{\nu}-\nu)^{2},\ell^{\prime\prime\prime}(\nu))$ $=\frac{1}{I^{2}}\sum_{i=1}^{n}cov((U^{i})^{2},W^{i})=O(n^{-1}).$
Therefore,
$IE(\hat{\nu}-\nu)+\frac{J}{I}+\frac{K}{2I}=0\quad\quad b:=E(\hat{\nu}-\nu)=\frac{1}{2I^{2}}(K+2J),$
which is of order $n^{-1}$. ∎
Appendix G Extra Results of robust regression simulations
G.1 Results when $p=3$
Figure 12: Overall RMSE of $\hat{\beta}$ with 10%, 20%, and 30% two-point contaminated errors when $p=3$
Refer to caption Refer to caption Refer to caption
Figure 13: Overall RMSE of $\hat{\beta}$ with 10% 20%, and 30% $t(2)$ contaminated errors when $p=3$
Refer to caption Refer to caption Refer to caption
G.2 Results when $p=80$
Figure 14: Overall RMSE of $\hat{\beta}$ with 10% 20%, and 30% $\chi^{2}(4)-4$ contaminated errors when $p=80$
Refer to caption Refer to caption Refer to caption
Figure 15: Overall RMSE of $\hat{\beta}$ with 10% 20%, and 30% $t(2)$ contaminated errors when $p=80$
Refer to caption Refer to caption Refer to caption
Appendix H Proof of Proper Posterior distribution for using $I_{\nu\nu}^{1/2}(\nu)$
We followed the proof by (Fonseca et al., 2008). Recall
$I_{\nu\nu}^{1/2}(\nu)=\left[\psi^{\prime}\left(\frac{\nu}{2}\right)-\psi^{\prime}\left(\frac{\nu+1}{2}\right)-\frac{2(\nu+5)}{\nu(\nu+1)(\nu+3)}\right]^{1/2}$
Corollary 1.1.
The $I_{\nu\nu}^{1/2}$ prior for $\nu$ is a continuous function in $0,\infty)$ and $I_{\nu\nu}^{1/2}=\mathcal{O}(\nu^{-1})$ as $\nu\rightarrow 0$ and $I_{\nu\nu}^{1/2}=\mathcal{O}(\nu^{-2})$ as $\nu\rightarrow\infty$.
Proof.
$\frac{2(\nu+5)}{\nu(\nu+1)(\nu+3)}=\mathcal{O}(\nu^{-2})$. Since $\psi^{\prime}(x)=\sum_{n=0}^{\infty}\frac{1}{(x+n)^{2}}$, as $\nu\rightarrow 0$, $\psi^{\prime}\left(\frac{\nu}{2}\right)=O(\nu^{-2})$. As $\nu\rightarrow 0$, by Stirling’s asymptotic formula, $\psi^{\prime}(x)\simeq x^{-1}+(2x^{2})^{-1}$ for large $x$. Therefore, we have
$I_{\nu\nu}\simeq=\frac{2\nu^{2}+6\nu+2}{\nu^{2}(\nu+1)^{2}}-\frac{2(\nu+5)}{\nu(\nu+1)(\nu+3)}=\frac{10\nu+6}{\nu^{2}(\nu+1)^{2}(\nu+3)}=\mathcal{O}(\nu^{-4}).$
Theorem 2.
Suppose $n\geq p+1$, $I_{\nu\nu}^{1/2}$ yields a proper posterior distribution.
Proof.
The results follows from Corollary 1.1 and the theorem 1 at page 156 of Fernández and Steel (1999). ∎